GFTYPE(1) GFTYPE(1) NAME gftype−translateagenericfontfileforhumanstoread SYNOPSIS gftype[−i][−m]gffilename[outputfilename] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thegftypeprogramtranslatesagf(genericfont)fileoutputby,forexample,mf(1),toafilethat humanscanread.Italsoservesasagffile-validatingprogram(i.e.,ifgftypecanreadit,it’scor-rect)andasanexampleofagf-readingprogramforothersoftwarethatwantstoreadgffiles. Thegffilenameonthecommandlinemustbecomplete.Becausetheresolutionispartofthe extension,itwouldnotmakesensetoappendadefaultextensionasisdonewithTEXorDVIreadingsoftware.Ifnooutputfilenameisspecified,theoutputgoestostdout. OPTIONS Theoutputfileincludesbitmapimagesofthecharacters,if−iisgivenonthecommandline,and asymboliclistingofthegfcommands,if−misspecified. ENVIRONMENT gftypelooksforgffilenameusingtheenvironmentvariableGFFONTS.Ifthatisnotset,it usesthevariableTEXFONTS.Ifthatisnotset,itusesthesystemdefault. SEEALSO dvitype(1),pktype(1). DonaldE.Knuthetal.,METAFONTware. AUTHORS DavidFuchsandDonKnuthwrotetheprogram.ItwaspublishedaspartoftheMETAFONTware technicalreport,availablefromtheTEXUsersGroup.PaulRichardsoriginallyportedittoUnix. Web2C7.32x 14December1993 1